In Cartesian coordinates, verify that ∇•(MA) = A •∇M + M∇•A where A = xyz(ax + ay + az) and
M= 3xy + 4zx by carrying out the indicated derivatives.
1
Expert's answer
2020-10-19T13:23:01-0400
"A \cdot∇M=(xyz(ax + ay + az))∇(3xy + 4zx)\\=(xyz(ax + ay + az))(3y+4z,3x,4x)"

"M \cdot∇A=(3xy + 4zx)∇(xyz(ax + ay + az))\\=a(3xy + 4zx)\cdot\\(yz(2x+y+z),xz(x+2y+z),yx(x+y+2z))"

"∇(MA)=∇(xyz(ax + ay + az)(3xy + 4zx))=a(3xy + 4zx)\cdot\\(yz(2x+y+z),xz(x+2y+z),yx(x+y+2z))+\\(xyz(ax + ay + az))(3y+4z,3x,4x)"

Thus,


"∇(MA)=A \cdot∇M+M \cdot∇A"